# Nicholas Michel    Mohamed E. Najd    Ghada Almashaqbeh

### v3

@@ 4.4 Security
Since $\mathsf{ammBoost}$ delegates the processing of the AMM transactions to the sidechain, and introduces pruning and state synchronization, we show that under this new architecture, the security and correct operation (i.e., safety and liveness) of the underlying AMM are preserved. In Appendix 0.A, we prove the following theorem:
Theorem 4.1
@@
5 Implementation
To assess the performance gains that $\mathsf{ammBoost}$ provides for AMMs, we implement a proof-of-concept and conduct various experiments. [^12] We will open source our implementation. We chose a Uniswap-inspired use case to represent the underlying AMM. This section discusses the implementation details, while the experimental setup and results are discussed in the next section.
Sidechain implementation. For the sidechain, we use the chainBoost implementation from [3]. This code uses cryptographic sortition-inspired election mechanism [48] for the committee election, and this committee runs a BLS collective signing (CoSi)-based PBFT-based consensus algorithm [7]. We add the modifications needed for $\mathsf{ammBoost}$; our sidechain has its own miners, and adopts the summary rules defined in Section 4. We also modify the syncing process to be an invocation to the $\mathsf{Sync}$ function in $\mathsf{TokenBank}$ authenticated using the threshold signature-based quorum certificate discussed earlier. We use a simplified version of the golang BLS library [68], and a pre-generated key to sign the $\mathsf{Sync}$ transaction. Furthermore, our sidechain implements two extra functions to aid in performing the AMM functionality:
  • $\mathsf{CreateTx}_{\mathsf{Sync}}$: creates the $\mathsf{Sync}$ call inputs based on the summary-block, including the payouts, the updated liquidity positions, and the pool liquidity balance. This is called by the sidechain committee leader at the end of each epoch.
[…]
@@ 5 Implementation
Use case: Uniswap-inspired AMM. We implement swaps, mints, burns, and collects using the same logic as in Uniswap (Appendix 0.B). We do not implement flashes since they represent a very small portion of the traffic, and thus will not impact the performance gains we report. For simplicity, our use case implementation manages a single pool. We deployed two standard ERC20 contracts to provide the token pair traded in this pool and used in both the ammBoost and baseline experiments. Naturally, to test Uniswap V3 in an isolated environment, we use the UniswapV3Factory contract to deploy a new liquidity pool which hosts the two ERC20 tokens. In order to test against the baseline implementation of Uniswap V3, we wrote and deployed a smart contract to interface with the various Uniswap contracts as detailed in the Uniswap documentation [26]. This interface contract routes swaps to the swapRouter contract and mints/burns/collects to the NFPM. Additionally, it manages all of the NFT liquidity positions (through the ERC721Reciever interface) created by the users in our experiments.
Traffic generation. Users generate traffic on both the mainchain and the sidechain. On the sidechain, the traffic follows the same distribution as in Uniswap (see Appendix 0.C), i.e., 93.19% of the traffic is $\mathsf{tx}_{\mathsf{swap}}$, 2.14% is $\mathsf{tx}_{\mathsf{mint}}$, 2.38% is $\mathsf{tx}_{\mathsf{burn}}$ and 2.27% is $\mathsf{tx}_{\mathsf{collect}}$. Our implementation provides configuration settings to modify the distribution and volume of the generated transactions to test their impact on the reported performance metrics.
6 Performance Evaluation
6.1 Experiment Setup
We deploy our system on a computing cluster composed of 8 hypervisors, each running a 12-Core, 130 GiB RAM, VM, connected with 1 Gbps network link. This setup is capable of running around 8000 sidechain miners. Unless stated otherwise, an experiment length is 11 epochs, each of which consists of 30 sidechain rounds (a round lasts 7 sec). Our default meta-block size is 1 MB and a sidechain committee contains 500 miners. We deploy 100 AMM users, generating traffic that arrives at a constant rate of $\rho=\lceil\frac{V_{D}\times b_{t}}{3600\times 24}\rceil$ where $V_{D}=25\times 10^{6}$ is the chosen daily volume of transactions.
In reporting the results, we measure the following metrics:
@@ 6.1 Experiment Setup
    Throughput: Number of transactions processed per second.
@@
    Sidechain transaction latency: The delay between a transaction submission and its appearance in a meta-block. [^14] To obtain an accurate representation of this metric, and thus process a comparable amount of traffic, we empty the transaction queues after the end of each run.
@@
6.2 Comparison with the Baseline
@@
On-chain (itemized) per-operation overhead. We evaluate the overhead of the deposit and the syncing processes in $\mathsf{ammBoost}$, and compare that to the baseline Uniswap on-chain operations. We set the daily volume $V_{D}$ to be $500$K transactions (10x Uniswap). We use a Gas Profiler [23] to measure the gas cost of the different components of the $\mathsf{Sync}$ transaction. As shown in Table 6.2, we find that storing the state of the liquidity positions is the most expensive, as each consists of 192 bytes (or 6 words), incurring 22,100 gas units per word. The same gas cost of 22,100 gas unit per word is incurred when storing the liquidity pool balance. Each payout transaction incurs a constant fee of 15,771 units. A deposit of two tokens incurs a total cost of 105,392 units. The threshold signature-based quorum certificate incurs a fixed fee that corresponds to the gas cost of the BN256 operations needed for verification, and a fee that is proportional to the length of the summary data structure.
Table 1: Mainchain latency and itemized gas cost for $\mathsf{ammBoost}$ operations ($|sum|$ is the size of the summaries).
@@
Table 2: Mainchain latency and gas cost for Uniswap.
@@
Table 3: Operation storage overhead.
@@
Overall, the gas cost of the $\mathsf{Sync}$ call is affected by the number of positions processed in an epoch, the number of deposits made for an epoch, and the updated pool liquidity balance; this cost does not scale with the number of processed transactions, but rather with the number of clients and liquidity providers. On the other hand, in baseline Uniswap the gas cost is proportional to the total generated traffic, where the numbers in Table 2 are per one transaction from each type. For the average latency, a $\mathsf{Sync}$ transaction does not depend on any other mainchain transactions, so it is confirmed within one block on average. However, since a two-token deposit depends on two ERC20 approvals, and performs 2 transfers, it takes around 4 blocks in our experiments. The same behavior is observed in our Uniswap baseline, as a swap requires 1 approval from the user and a mint requires 2 approvals. Since a two-token deposit depends on two ERC20 approvals, and performs two transfers, it will take at least four rounds if the operations are done sequentially. For our Uniswap baseline, as a swap requires one approval, it takes a minimum of two rounds to be processed; a mint takes three rounds at least, at it requires two approvals.
We also report the per-operation storage cost. In particular, we report the cost breakdown for the $\mathsf{Sync}$ call on the mainchain and the summary-block size for $\mathsf{ammBoost}$, and the transaction sizes for baseline Uniswap on the mainchain. For the $\mathsf{Sync}$ call, the sizes of payout and position entries vary greatly between the summarized changes in a summary-block and the $\mathsf{Sync}$ inputs submitted to the mainchain. This is due to the difference in encoding and binary packing between the side and main chains. On the mainchain, Ethereum’s application binary interface (ABI) packing keeps track of the data and all the information needed to reinterpret it back, while on the sidechain we use simple binary packing. We also have an extra 6 words (192 bytes) storage overhead on the mainchain needed for the BLS signature and its public key (namely, $\mathsf{vk}_{c}$) for authenticating the $\mathsf{Sync}$ call. We present our findings in Table 3.
As shown, for Uniswap, we notice that the transactions on Sepolia are smaller than the ones we observe on Ethereum (Appendix 0.C). This is because these chains use different Uniswap transaction routers. The calls to the universal router used on Ethereum end up requiring more arguments, resulting in longer transactions. Uniswap on Sepolia deploys a simpler transaction router. Of note is that the simple router contract (Uniswap V3 router) and the more complex of the two (the Universal router) are both available on Ethereum, but the Universal router is not deployed to Sepolia.
@@
In general, and as will discussed next, Uniswap incurs a larger storage cost as its transactions are quite large and all are logged on the mainchain. While for $\mathsf{ammBoost}$, only (the less frequent) $\mathsf{Sync}$ call transaction is logged on the mainchain.
@@
As shown in Figure 6, even if the sync transactions end up being heavy on gas as the number of positions and payouts increases, we achieve a 96.05% gas reduction when compared to Uniswap Sepolia. The high gas cost of the $\mathsf{Sync}$ transaction is offset by it being uncommon (one occurrence per epoch). On the other hand, the gas cost of swaps, mints, burns, and collects in Uniswap are high since all are processed on the mainchain (while in $\mathsf{ammBoost}$ these are processed on the sidechain). A similar trend is observed for the mainchain state growth, where $\mathsf{ammBoost}$ provides 93.42% decrease in growth compared to Uniswap on Sepolia, and 97.60% decrease when compared to Uniswap on production Ethereum. [^15] The growth for Uniswap on production Ethereum is calculated by multiplying the count of each transaction type in our experiment by its size as reported in Appendix 0.C
6.3 Impact of Parameter Configuration
We study the impact of parameter configuration on $\mathsf{ammBoost}$’s performance, including traffic amount and distribution, block size, sidechain round duration, and number of rounds per epoch.
Scalability. In this experiment, we test the scalability of $\mathsf{ammBoost}$ (for a single pool) to understand its behavior under heavy traffic. We follow the same traffic distribution as in Uniswap and vary the daily volume $V_{D}\in\{50K,500K,5M,25M\}$. We record the impact on throughput and transaction/payout latency as shown in Table 4.
Table 4: Scalability of $\mathsf{ammBoost}$.

### v4

@@ IV-D Security
Since $\mathsf{ammBoost}$ delegates AMM transaction processing to the sidechain, and introduces pruning and state synchronization, we show that under this new architecture, operation correctness and security (i.e., safety and liveness) of the underlying AMM are preserved. In Appendix B, we prove the following theorem:
Theorem 1.
@@
Briefly, $\mathsf{ammBoost}$ preserves the correct behavior of the AMM since it processes the sidechain workload using the same logic adopted by the AMM itself. Thus, all transaction types—swaps, mints, collects, and burns—will produce the same outcome, or state changes, as if they are processed on the mainchain. Also, since the sidechain adopts a secure PBFT consensus protocol, that satisfies safety and liveness, the committee only agrees on valid meta-blocks that contain transactions and state changes conforming with the AMM operation and rules. Also, they only agree on valid summary-blocks produced based on the summary rules set in the system that correctly reflect these state changes. Furthermore, liveness of this protocol guarantees that valid submitted transaction will be processed/published on this sidechain. The sidechain also satisfies public verfiability since meta-blocks do get pruned until their sync-transaction is confirmed on the mainchain.
For the AMM base smart contract, recall that it is a dApp deployed on top of a secure smart contract-enabled blockchain (i.e., the mainchain), and its safety/liveness is guaranteed by the security of this chain. That is, mainchain miners will process only valid transactions (including sync-transactions authenticated using TSQC) based on $\mathsf{TokenBank}$ code, and any invalid transaction will be rejected. Lastly, we note that $\mathsf{ammBoost}$ handles any interruptions on the sidechain side and roll-backs on the mainchain, as discussed in the previous section, using the leader change and mass-syncing mechanisms.
V Implementation
To assess the performance gains of $\mathsf{ammBoost}$, we implement a proof-of-concept and conduct various experiments (our code can be found at [86]). We chose a Uniswap-inspired use case to represent the underlying AMM. This section discusses the implementation, while the next section shows our results.
Sidechain implementation. We use the code from [87], which employs cryptographic sortition-inspired election mechanism [76] for committee election, and this committee runs a BLS collective signing (CoSi)-based PBFT-based consensus [88]. We add the modifications needed for $\mathsf{ammBoost}$; our sidechain has its own miners, and adopts the summary rules from Section IV. We also modify the syncing process to be an invocation to the $\mathsf{Sync}$ function in $\mathsf{TokenBank}$ authenticated using TSQC for which we use a simplified version of the golang BLS library [89], and a pre-generated key to sign the $\mathsf{Sync}$ transaction. Furthermore, our sidechain implements two extra functions to aid in performing the AMM functionality:
  • $\mathsf{CreateTx}_{\mathsf{Sync}}$: invoked by the sidechain committee at the end of any epoch to create the $\mathsf{Sync}$ call inputs based on the summary-block, including payouts, updated liquidity positions, and liquidity pool balance.
  • $\mathsf{SnapshotBank}$: retrieves users’ deposits at the beginning of an epoch. $\mathsf{ammBoost}$ retrieves pool balances only for newly created pools, their updated balances can be computed by the sidechain based on the traffic it processes.
@@ V Implementation
Use case: Uniswap-inspired AMM. We implement swaps, mints, burns, and collects using the same logic as in Uniswap (Appendix C). We do not implement flashes since they represent a very small portion of the traffic, and thus will not impact the performance gains we report. For simplicity, our implementation manages a single pool, where we deployed two standard ERC20 contracts to provide the token pair traded in this pool and used in both the ammBoost and baseline experiments. To test Uniswap V3 in an isolated environment, we use the UniswapV3Factory contract to deploy a pool hosting the two ERC20 tokens. To test against the baseline implementation of Uniswap V3, we wrote and deployed a smart contract to interface with the various Uniswap contracts as detailed in the Uniswap documentation [96]. This interface contract routes swaps to the swapRouter contract and mints/burns/collects to the Nonfungible position manager (NFPM) contract. Additionally, it manages all liquidity positions (through the ERC721Reciever interface) created by the users in our experiments.
Traffic generation. Users generate traffic on both the mainchain and the sidechain. On the sidechain, the traffic follows the same distribution as in Uniswap (see Appendix D), i.e., 93.19% of the traffic is $\mathsf{tx}_{\mathsf{swap}}$, 2.14% is $\mathsf{tx}_{\mathsf{mint}}$, 2.38% is $\mathsf{tx}_{\mathsf{burn}}$ and 2.27% is $\mathsf{tx}_{\mathsf{collect}}$. Our implementation provides configuration settings to modify the distribution and volume of the generated transactions to test their impact on the reported performance metrics.
VI Performance Evaluation
VI-A Experiment Setup
We deploy our system on a computing cluster composed of 8 hypervisors, each running a 12-Core, 130 GiB RAM, VM, connected with 1 Gbps network link. This setup is capable of running around 8000 sidechain miners. Unless stated otherwise, an experiment length is 11 epochs, each of which consists of 30 sidechain rounds (a round lasts 7 sec). [^5] This is a reasonable round duration in light of recent systems, e.g., in Polygon [97] and Algorand [98] a round lasts for 2 and 2.8 sec, respectively. Our default meta-block size is 1 MB and a sidechain committee contains 500 miners—a comparable size to those in [12, 13, 10] (we test the impact of varying the committee size in Appendix E). We deploy 100 AMM users, generating traffic that arrives at a constant rate of $\rho=\lceil\frac{V_{D}\times b_{t}}{3600\times 24}\rceil$ where $V_{D}=25\times 10^{6}$ is the chosen daily volume of transactions. We measure the following metrics:
@@ VI-A Experiment Setup
    Throughput: Number of transactions processed per sec.
@@
    Sidechain transaction latency: The delay between a transaction submission and its appearance in a meta-block. For accurate latency representation, and thus processing a comparable traffic amount, we empty the transaction queues after the end of each run.
@@
VI-B Comparison with the Baseline
@@
On-chain (itemized) per-operation overhead. We evaluate the overhead of the deposit and the syncing processes in $\mathsf{ammBoost}$, and compare that to the baseline Uniswap on-chain operations. We set the daily volume $V_{D}$ to be $500$K transactions (10x Uniswap). We use a Gas Profiler [99] to measure the gas cost of the different components of the $\mathsf{Sync}$ transaction. As shown in Table VI-B, we find that storing the state of the liquidity positions is the most expensive, as each consists of 192 bytes (or 6 words), incurring 22,100 gas units per word. The same gas cost of 22,100 gas unit per word is incurred when storing the liquidity pool balance. Each payout transaction incurs a constant fee of 15,771 units. A deposit of two tokens incurs a total cost of 105,392 units. The threshold signature-based quorum certificate incurs a fixed fee that corresponds to the gas cost of the BN256 operations needed for verification, and a fee that is proportional to the length of the summary data structure.
TABLE II: Mainchain latency and itemized gas cost for $\mathsf{ammBoost}$ operations ($|sum|$ is the size of the summaries).
@@
TABLE III: Mainchain latency and gas cost for Uniswap.
@@
Overall, the gas cost of the $\mathsf{Sync}$ call is affected by the number of positions processed in an epoch and the updated pool balance; this cost does not scale with the number of processed transactions, but rather with the number of clients and liquidity providers. On the other hand, in baseline Uniswap the gas cost is proportional to the total generated traffic, where the numbers in Table III are per one transaction from each type. For the average latency, a $\mathsf{Sync}$ transaction does not depend on any other mainchain transactions, so it is confirmed within one block on average. However, since a two-token deposit depends on two ERC20 approvals, and performs 2 transfers, it takes around 4 blocks in our experiments. The same behavior is observed in the Uniswap baseline, as a swap requires one approval from the user, it takes a minimum of two rounds to be processed, and a mint requires two approvals, thus it takes three rounds at least, if the operations are done sequentially.
We also report the per-operation storage cost. In particular, we report the cost breakdown for the $\mathsf{Sync}$ call on the mainchain and the summary-block size for $\mathsf{ammBoost}$, and the transaction sizes for baseline Uniswap on the mainchain. For the $\mathsf{Sync}$ call, the sizes of payout and position entries vary greatly between the summarized changes in a summary-block and the $\mathsf{Sync}$ inputs submitted to the mainchain. This is due to the difference in encoding and binary packing between the side and main chains. On the mainchain, Ethereum’s application binary interface (ABI) packing keeps track of the data and all the information needed to reinterpret it back, while on the sidechain we use simple binary packing. We also have an extra 192 bytes storage overhead on the mainchain needed for the BLS signature and its public key (namely, $\mathsf{vk}_{c}$) for authenticating the $\mathsf{Sync}$ call. Our findings are in Table IV.
For Uniswap, we notice that the transactions on Sepolia are smaller than the ones we observe on Ethereum (Appendix D). This is because these chains use different Uniswap transaction routers. The calls to the universal router used on Ethereum end up requiring more arguments, resulting in longer transactions. Uniswap on Sepolia deploys a simpler transaction router and without the universal router (while Uniswap V3 on Ethereum deploys both contracts).
In general, Uniswap incurs a larger storage cost as its transactions are quite large and all are logged on the mainchain. While for $\mathsf{ammBoost}$, only (the less frequent) $\mathsf{Sync}$ call transaction is logged on the mainchain.
TABLE IV: Operation storage overhead.
@@
As shown in Figure 5, even if the sync transactions end up being heavy on gas as the number of positions and payouts increases, we achieve a 96.05% gas reduction when compared to Uniswap Sepolia. The high gas cost of the $\mathsf{Sync}$ transaction is offset by it being uncommon (one occurrence per epoch). On the other hand, the gas cost of swaps, mints, burns, and collects in Uniswap are high since all are processed on the mainchain (while in $\mathsf{ammBoost}$ these are processed on the sidechain). A similar trend is observed for the mainchain state growth, where $\mathsf{ammBoost}$ provides 93.42% decrease in growth compared to Uniswap on Sepolia, and 97.60% decrease when compared to Uniswap on production Ethereum. [^6] The growth for Uniswap on production Ethereum is calculated by multiplying the count of each transaction type in our experiment by its size as reported in Appendix D.
VI-C Impact of Parameter Configuration
We study the impact of parameter configuration on $\mathsf{ammBoost}$’s performance, including scalability, traffic distribution, block size, sidechain round duration, number of rounds per epoch, and committee size. Due to space limitation, we only show the full results for the scalability experiment, while the rest can be found in Appendix E.
Scalability. In this experiment, we test the scalability of $\mathsf{ammBoost}$ (for a single pool) to understand its behavior under heavy traffic. We follow the same traffic distribution as in Uniswap and vary the daily volume $V_{D}\in\{50K,500K,5M,25M\}$. We record the impact on throughput and transaction/payout latency as shown in Table V.
Latency-wise, we achieve a quasi-instant trend when the daily volume is 50K-5M (transactions that arrive at the beginning of the round get processed within the same round, while the residual amount of latency is due to transactions generated close to the epoch end and processed in the next epoch). This leads to payouts being processed within one epoch, with the transactions happening earlier in the epoch having a higher payout latency than the ones happening towards the end, and averaging at half an epoch plus the time needed to confirm the $\mathsf{Sync}$ transaction on the mainchain $(\sim 120s)$. Transaction congestion happens when the daily volume is 25M, resulting in higher average latency as the table shows.
TABLE V: Scalability of $\mathsf{ammBoost}$.